2532dj:

An observer (O) is located 660 feet from a tree (T). The observer notices a hawk (H) flying at a 35° angle of elevation from his line of sight. How high is the hawk flying over the tree? You must show all work and calculations to receive full credit.

surjithayer:

 \[\frac{ h }{ 600 }=\sin ~35\] h=660 sin 35 \[h \approx 378.56 ~ft\]

surjithayer:

correction in fig it is 660 ft and not 600

surjithayer:

Created with Raphaël35600hReply Using Drawing \[\frac{ h }{ 660 }=\sin ~35\] h=660 sin 35 \[h \approx 378.56 ~ft\]